# Open Hardware in Quantum Technology

### v1

@@ II.2.4 Optimal control, calibration and characterization
Example 5. OPTIMAL CONTROL, CALIBRATION AND CHARACTERIZATION.
C3: An integrated tool-set for Control, Calibration and Characterization.
The C3-Toolset package provides an API for directly interfacing with hardware to use the pulses generated by optimal control and further optimize them using closed-loop calibration algorithms. Users have access to a high-level Qiskit interface that allows them to run gate or pulse level quantum circuits on the full-physics noisy high-fidelity differentiable simulator. In recent years, a variety of machine learning for quantum control techniques have been closely integrated in the quantum device bring up process. Besides the previously mentioned process of learning model parameters, reinforcement learning (RL) has been particularly useful in a variety of applications. The open-source library rbqoc [123] details a technique for noise-robust control using RL while C3-Toolset uses RL for both learned optimizers and Bayesian experiment design.
@@ II.4.2 Deeply collaboratively testbeds
While cloud-based platforms allow scientists and the public alike to submit their circuits and receive the results, deeper, customized access to the full stack to enable more involved experiments and R&D efforts is generally not possible on these platforms. To provide such access to users in Academia, Industry, and National Labs, the US Department of Energy has funded two testbed programs: QSCOUT (Quantum Scientific Computing Open User Testbed), based on trapped ions and located at Sandia National Laboratories, and AQT (Advanced Quantum Testbed), based on superconducting qubits and located at Lawrence Berkeley National Laboratory. A similar network called Qu-Test [143] managed by TNO at Delft, Netherlands has recently been kicked off with the goal of providing a federated network of testbeds by bringing together 13 service providers and 11 industrial users from the European quantum community. These platforms allow low-level access to the full stack of quantum hardware (including programming languages [144, 145], gate-pulse shaping [103], noise injection, unique qubit/qudit states, specialized calibrations, comparison of compilation techniques, etc.) allowing users to probe how the machines actually work and how to make them work better. The testbeds foster deep collaborations between the host laboratory and the users.
@@ III Discussion: Current Gaps & Future Recommendations
Support of OQH in education and academia. We believe it is important to foster OSS and OQH adoption in education and research. One actionable item is to encourage schools and universities to include in course material and labs the usage of open-source software [158] and open hardware [159, 160, 161]. Programs that prepare the future quantum workforce are fundamental to broaden the acquisition of talents [24, 5, 6]. This includes developing communities across domain knowledge.
@@ V Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Accelerated Research in Quantum Computing under Award Number DE-SC0020266. AM acknowledges support from the PNRR MUR project PE0000023-NQSTI.

### v2

@@ II.2.4 Optimal control, calibration and characterization
Example 5, Control & Data Acquisition: “Qudi: A modular python suite for experiment control and data processing.” Qudi provides a flexible and modular framework for controlling and analysing data from confocal microscopy, quantum optics and quantum information experiments primarily based on color centers in diamond. Developed using Python and Qt, it is a feature rich open-source solution providing a GUI, real time data visualization, distributed execution over networks, Jupyter notebook interface, configuration management, and data recording [58]. A key feature of Qudi is its modular design consisting of functionally independent parts that are orchestrated by a manager component with a declarative configuration management system. This also makes it very easy to extend the framework to support new devices and experiments. The entry point for any user is the start.py file which loads the core components that take care of configuration parsing, module management, error logging, and remote network access. There is a three-tier architecture consisting of logic modules, hardware modules and GUI modules ensuring a clear separation of concern and allowing independent use, development and deployment.
@@
Example 6. Optimal Control, Calibration and Characterization:
“C3: An integrated tool-set for Control, Calibration and Characterization.”
The C3-Toolset package provides an API for directly interfacing with hardware to use the pulses generated by optimal control and further optimize them using closed-loop calibration algorithms. Users have access to a high-level Qiskit interface that allows them to run gate or pulse level quantum circuits on the full-physics noisy high-fidelity differentiable simulator. Libraries like QuOCS [127] also provide an extensive suite of model-based optimal control and pulse-shaping for a wide variety of hardware platforms. In recent years, a variety of machine learning for quantum control techniques have been closely integrated in the quantum device bring up process. Besides the previously mentioned process of learning model parameters, reinforcement learning (RL) has been particularly useful in a variety of applications. The open-source library rbqoc [128] details a technique for noise-robust control using RL.
@@ II.4.2 Deeply collaboratively testbeds
While cloud-based platforms allow scientists and the public alike to submit their circuits and receive the results, deeper, customized access to the full stack to enable more involved experiments and R&D efforts is generally not possible on these platforms. To provide such access to users in Academia, Industry, and National Labs, the US Department of Energy has funded two testbed programs: QSCOUT (Quantum Scientific Computing Open User Testbed), based on trapped ions and located at Sandia National Laboratories, and AQT (Advanced Quantum Testbed), based on superconducting qubits and located at Lawrence Berkeley National Laboratory. A similar network called Qu-Test [148] managed by TNO at Delft, Netherlands has recently been kicked off with the goal of providing a federated network of testbeds by bringing together 13 service providers and 11 industrial users from the European quantum community. Likewise, in the UK, the National Quantum Computing Centre (NQCC) is building, testing and hosting full stack quantum computing solutions [65]. Through close collaboration between industry, government and academia, the NQCC is initially focusing on developing demonstrator platforms which provide direct access to quantum computing resources while helping to also develop a robust supply chain. A recently concluded funding call by the Innovate UK [149] earmarked 30 million GBP to build, commission, and validate an operational quantum computing testbed in the UK as part of the NQCC.
In Canada, the Open Quantum Design (OQD)initiative aims at building a completely open-source and open-hardware quantum computer based on ions. OQD has taken the form of a full-fledged non-profit organization arising from previous experimental and theoretical work performed at the Institute for Quantum Computing and Perimeter Institute.
These platforms allow low-level access to the full stack of quantum hardware (including programming languages [150, 151], gate-pulse shaping [106], noise injection, unique qubit/qudit states, specialized calibrations, comparison of compilation techniques, etc.) allowing users to probe how the machines actually work and how to make them work better. The testbeds foster deep collaborations between the host laboratory and the users.
@@ III Discussion: Current Gaps & Future Recommendations
Support of OQH in education and academia. We believe it is important to foster OSS and OQH adoption in education and research. One actionable item is to encourage schools and universities to include in course material and labs the usage of open-source software [164] and open hardware [165, 166, 167]. Programs that prepare the future quantum workforce are fundamental to broaden the acquisition of talents and develop new career paths [24, 5, 6]. This includes developing communities across domain knowledge.
@@ V Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Accelerated Research in Quantum Computing under Award Number DE-SC0020266. AM acknowledges support from the PNRR MUR project PE0000023-NQSTI. QubiC and QUASAR research was funded by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research Quantum Testbed Program under contract DE-AC02-05CH11231